40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 19 04:25:54 2011, Kevin, Update, Electronics, POX Transfer Functions 
    Reply  Wed Jan 19 10:45:00 2011, Koji, Update, Electronics, POX Transfer Functions 
       Reply  Wed Jan 19 17:00:23 2011, Kevin, Update, Electronics, POX Transfer Functions 
          Reply  Thu Jan 20 01:50:30 2011, Kevin, Update, Electronics, POX Transfer Functions 
Message ID: 4170     Entry time: Wed Jan 19 17:00:23 2011     In reply to: 4169     Reply to this: 4172
Author: Kevin 
Type: Update 
Category: Electronics 
Subject: POX Transfer Functions 

 The value of I_dc was a mistake. The value should be 240 µA.

The widths of the resonance peaks are listed below the fits to each peak on the wiki.

ELOG V3.1.3-